Circular No. 7426

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 J4 (SOHO) D. A. Biesecker, Emergent Information Technologies, Inc., and God-

dard Space Flight Center, reports observations of another Kreutz sungrazing comet found by M. Oates and X. Leprette in LASCO C2 data at the SOHO website. The comet is also visible in C3 data, and a very short tail was visible. Astrometry (measured by D. Hammer and Biesecker, reduced by B. G. Marsden) and orbital elements appear on MPEC 2000-K08.

2000 UT  α_{2000}  δ_{2000}
May 14.904  3^{h}32.^{m}9  +16^{o}58′

CI AQUILAE J. C. Wilson, Cornell University; and K. R. Dunscombe, Palomar Ob-

servatory, report 0.75–2.5-μm spectroscopy of the probable nova in Aquila (IAUC 7409) using the Palomar 1.52-m telescope (+ Cornell/Massachusetts Slit Spectrograph) on May 9.493 UT (∼ 11 days after outburst): “Spectra were taken of the infrared bright object found at α = 18^{h}52^{m}04^{s}.0, δ = −1^{o}28′40′′ (equinox 2000.0), which correlates roughly with the suspected 1917 nova and eclipsing binary CI Aql. The spectrum shows nearly 30 emission lines, including Paschen- and Brackett-series hydrogen lines as well as many N I and O I features. He I (1.0830 μm) is self-absorbed. Hydrogen line ratios differ profoundly from the recombination values. Lyβ fluoresced O I lines (0.8446 and 1.1287 μm) are very strong. The K magnitude is estimated to be 7–8.”

M. O. Jesacher, S. J. Kautsch, S. Kimeswenger, M. S. Mühlbacher, W. Saurer, S. Schmeja, and C. K. Scholz, University of Innsbruck, report BVR CCD magnitudes (± 0.05 mag) obtained with the Innsbruck 0.60m telescope on May 10.97 UT and infrared magnitudes obtained with the European Southern Observatory 1-m telescope (+ DENIS) on May 11.38: B = 10.49, V = 9.67, R = 8.48, I_{c} = 8.28, J = 6.35, K_{s} = 6.12.

V4334 SAGITTARII Kimeswenger and Schmeja also report infrared-imaging photometry of

V4334 Sgr on May 6.292 UT at the European Southern Observatory 1-m telescope (+ DENIS): “The K flux has increased compared to recent reports (IAUC 7266, 7422), while the star faded in J and vanished in I (K = 7.03, J = 14.22, I > 19.1). Such a behavior cannot be described by cooling of the dust, but implies a strong further increase of the internal extinction.”

2000 May 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green